### v2

# DM-Bench: Benchmarking LLMs for Personalized Decision Making in Diabetes Management

Abstract
We present DM-Bench, the first benchmark designed to evaluate large language model (LLM) performance across real-world decision-making tasks faced by individuals managing diabetes in their daily lives. Unlike prior health benchmarks that are either generic, clinician-facing or focused on clinical tasks (e.g., diagnosis, triage), DM-Bench introduces a comprehensive evaluation framework tailored to the unique challenges of prototyping patient-facing AI solutions in diabetes, glucose management, metabolic health and related domains. Our benchmark encompasses 7 distinct task categories, reflecting the breadth of real-world questions individuals with diabetes ask, including basic glucose interpretation, educational queries, behavioral associations, advanced decision making and long term planning. Towards this end, we compile a rich dataset comprising one month of time-series data encompassing glucose traces and metrics from continuous glucose monitors (CGMs) and behavioral logs (e.g., eating and activity patterns) from 15,000 individuals across three different diabetes populations (type 1, type 2, prediabetes/general health and wellness). Using this data, we generate a total of 360,600 personalized, contextual questions across the 7 tasks. We evaluate model performance on these tasks across 5 metrics: accuracy, groundedness, safety, clarity and actionability. Our analysis of 8 recent LLMs reveals substantial variability across tasks and metrics; no single model consistently outperforms others across all dimensions. By establishing this benchmark, we aim to advance the reliability, safety, effectiveness and practical utility of AI solutions in diabetes care.
1 Introduction
Individuals living with diabetes must continuously manage their blood glucose levels to avoid adverse health consequences, a process that involves frequent, complex decision-making. This decision making process is highly personalized and context-dependent, varying between individuals and across diabetes populations. For example, individuals with type 1 diabetes often focus on insulin titration and maintaining glucose within a tight range, while those with type 2 diabetes who are not on insulin may prioritize reducing glycemic variability and achieving broader lifestyle goals such as weight loss. Diabetes management is increasingly supported by wearable devices including continuous glucose monitors (CGMs), which provide real-time glucose data, and other wearables like smart watches, smart rings and companion apps that allow users to log meals, track physical activity, and monitor behavioral patterns (Jafleh et al., 2024). These devices generate highly-granular longitudinal streams of personal health data over weeks, months and even years.
The explosion of rich personal health data presents a significant opportunity for Artificial Intelligence (AI) and particularly large language models (LLMs) to support individuals in managing their diabetes (Mahajan et al., 2025). In fact, exciting recent developments in both academia and industry have begun to explore the integration of LLMs into diabetes management contexts including for nutrition and glucose monitoring (Guan et al., 2023), answering medical questions (Hussain & Grundy, 2025), and generating insights and logging meals (Dexcom, 2024; 2025). As LLM capabilities continue to advance, especially in processing multimodal data and handling long, complex time-series, they offer immense potential for creating seamless patient-facing tools that deliver nuanced, actionable, context-aware and personalized insights and guidance, optimally leveraging the highly granular and longitudinal data generated by these wearable devices.
Despite this promise, fully realizing these benefits requires that AI models be developed and evaluated in safe, effective and trustworthy ways. A critical component is the establishment of robust, standardized benchmarks to guide model development, assess performance in real-world settings, and support transparent comparisons across models. Currently, there are no publicly available benchmarks designed to evaluate models on patient-facing decision-making tasks related to diabetes and glucose management. Recent efforts have focused on general purpose health benchmarks such as HealthBench (Arora et al., 2025), MedHELM (Bedi et al., 2025), MedCalc-Bench (Khandekar et al., 2024), and MedGPTEval (Xu et al., 2024), as well as benchmarks for electronic health records, e.g., EHRShot (Wornow et al., 2023) and EHRNoteQA (Kweon et al., 2024). While valuable, these benchmarks are not tailored to the unique needs of individuals managing diabetes. The few benchmarks that do focus on diabetes domains are clinician-facing, targeting tasks such as diagnosis, triage, and report summarization (Wei et al., 2024; Healey & Kohane, 2024; Healey et al., 2025). These efforts often involve small cohort sizes and fail to capture the nuanced, personalized, and context-dependent decision-making that individuals with diabetes engage in daily.
Figure 1: DM-Bench spans 7 real-world tasks capturing realistic user needs in diabetes management.
Therefore, we introduce DM-Bench  [^1] DM stands for Diabetes Management., a comprehensive evaluation framework tailored to the unique requirements of prototyping in the diabetes, glucose management and metabolic health domains. DM-Bench is the first large-scale LLM benchmark designed to evaluate model performance on real-world, patient-facing diabetes management tasks. Our benchmark spans 7 distinct task categories (see Figure 1), designed to encompass the breadth of decision-making questions individuals with diabetes ask. These range from basic glucose interpretation (“What is my time in range today?”), and behavioral associations (“Why did this salad cause a glucose spike?”), to decision making and planning (“What workouts from this past month consistently lower my glucose levels?”). We compile a rich dataset of one month of time-series CGM and behavioral data from 15,000 individuals across three populations: type 1 diabetes, type 2 diabetes, and prediabetes/general health and wellness. Using this data, we generate 360,600 personalized, contextual questions across the 7 task categories. To evaluate model performance, we develop multi-dimensional evaluation criteria for each task, covering 5 important metrics: accuracy, groundedness, safety, clarity, and actionability. Finally, we evaluate a diverse set of LLMs and find that no model consistently outperforms across all tasks and metrics, highlighting the need for continued improvement in LLMs for diabetes management.
We present the following contributions: (1) We develop DM-Bench, a novel benchmark to evaluate LLMs on patient-facing diabetes management tasks created from wearable device data from 15,000 users across 3 diabetes populations. We generate 360,600 personalized, contextual questions, covering 7 real-world diabetes management tasks. (2) We develop a multi-dimensional evaluation framework for each task crafted by domain experts based on 5 key metrics: accuracy, groundedness, safety, clarity, and actionability. (3) We present comprehensive evaluations of 8 open-source and proprietary LLM models of differing sizes, purposes, and model families using DM-Bench. By establishing this benchmark, we aim to advance the reliability, safety and effectiveness of LLMs in metabolic health and diabetes, ultimately driving meaningful improvements for those living with diabetes. While focused on diabetes management, this framework is extensible to other domains involving wearable devices and continuous monitoring, including preventative care, fitness optimization and the management of other chronic conditions, e.g., hypertension, obesity, and sleep disorders. This benchmark also provides a foundation for evaluating LLMs on contextual reasoning tasks using complex, longitudinal time-series data across broader health and wellness applications.
2 DM-Bench
Figure 2: DM-Bench overview.
In this section we present DM-Bench, a benchmark for evaluating LLMs on diabetes management decision making tasks. Figure 2 shows an overview and we walk through each component next.
2.1 Task Curation
To support user-facing decision making in diabetes management, we curated tasks representative of realistic, personalized questions and concerns that individuals with diabetes would ask an AI system (Figure 2a). Our goal was to cover a broad range of patient-facing scenarios while spanning different levels of task complexity. We defined three categories: Glucose Literacy & Trends; Behavior–Glucose Dynamics; and Personalized Decision Support, comprised of a total of 7 tasks (Figure 1). First, individuals newly diagnosed with diabetes may want to build foundational knowledge about diabetes: what it is, how it works, and how it affects their daily lives. This includes understanding diabetes-specific metrics such as glucose variability and time in range, as well as interpreting their own data (Task 1: Glucose Math). It also involves learning how diabetes influences lifestyle choices and routines (Task 2: Education). Second, individuals often want to understand how specific behaviors influence their glucose levels. This includes reasoning about immediate, simple associations, such as the effect of a single meal or a night of poor sleep (Task 3: Simple Reasoning), as well as more complex, longer-term interactions between multiple behaviors and glucose outcomes (Task 4: Advanced Reasoning). Finally, individuals with diabetes may seek support for future-oriented decisions. This includes making momentary choices (Task 5: Decision Making), developing structured plans (Task 6: Planning), and identifying concerning trends that may warrant medical attention (Task 7: Alert/Triage). An overview of each task is in Table 1.
2.2 Data Curation
To ensure DM-Bench reflects the diverse needs of real-world diabetes populations, we curated data from 15,000 individuals evenly distributed across three cohorts: prediabetes/health and wellness (HW), type 1 diabetes (T1D), and type 2 diabetes (T2D) (Figure 2b). Each user contributed 30 consecutive days of CGM data in mg/dL, recorded at 5-minute intervals. Data was collected between January and June 2025 and paired with time-aligned behavioral logs of meals, exercise, sleep, and other glucose metrics, as well as daily activity summaries (e.g., step count, average heart rate). Depending on the task, data were segmented into 1-day, 7-day, or 30-day windows. The most complete data from the 30 days were used for the 1-day and 7-day windows, meaning the consecutive days with the richest self-reported behavioral data across categories. We also generated user data from GlucoSynth, which provides highly realistic, differentially-private synthetic glucose traces (Lamp et al., 2023). Additional details, including LLM input formatting, are in Appendix A.3.
Table 1: Task overview including data used and question generation process.
Task Description Data Used Data Length Question Generation Process
1 Glucose Math Glucose and time 1 day 9 templated questions
2 Education - - LLM generated 100 questions per cohort and age group
3 Simple Reasoning Glucose, time, and behavior 1 day LLM generated 3 questions per user
4 Advanced Reasoning Glucose, time, and behavior 30 days LLM generated 3 questions per user
5 Decision Making Glucose, time, and behavior 7 days LLM generated 3 questions per user
6 Planning Glucose, time, and behavior 30 days LLM generated 3 questions per user
7 Alert/Triage Glucose, time, and behavior 30 days LLM generated 3 questions per user
$\dagger$ Synthetic glucose data from Glucosynth used (Lamp et al., 2023).
2.3 Question Generation
We generated personalized questions by combining user context with task-specific goals (Figure 2c). Generally, for most tasks (Figure 2c.1), an LLM receives a structured, task-specific prompt and user data (Figure 2c.11). The prompt instructs the model to generate 3 customized questions reflecting the user’s context, including their data and diabetes type, across 3 behavior domains: sleep, exercise, and meals, each of which directly influences glucose regulation and diabetes management (ADA, 2025). Each question should reference a different behavior domain, but behavior types can be repeated if data for a given behavior are missing, and if no behavior data is available then questions may instead focus on user glucose trends. To ensure high quality questions are generated, each question is then evaluated by an LLM evaluator (Figure 2c.12) across five binary metrics: fluency, relevance, originality, difficulty, and answerability (see Table 5a in the Appendix for additional details). Questions failing any metric are iteratively refined until achieving a perfect score (5/5) or reaching five attempts. We then perform a cross-check for originality across all questions generated for the same user, with additional refinement if questions are too similar (Figure 2c.13). Finally, a human expert reviewer manually confirms the quality of a sample of the questions. All LLM-based generation used Gemini 2.5 Flash configured with 0 thinking.
Specifically, for Task 1 (Glucose Math) we designed 9 question templates, with placeholders (e.g., [metric], [time period]) that are filled with variable options, such as time in range, variance, or specific time windows, customized to each user (see Table 7 in the Appendix). Questions span general trends as well as domain-specific measures, resulting in 135,000 total questions (9 per user). Task 2 (Education) focuses on conversational learning without user data, for which we generate 600 questions across age groups (adult, adolescent) and diabetes types (HW, T1D, T2D). Importantly, we represent both adult and adolescent age groups since there is an increasing prevalence of diabetes in children (CDC, 2024) and AI health benchmarks tend to overlook adolescents (Muralidharan et al., 2024). Tasks 3–7 use user data from 1-day (Task 3), 7-day (Task 5), or 30-day windows (Tasks 4, 6, 7) to generate 3 behavior-grounded questions per user (45,000 per task). This framework results in 360,600 diverse, personalized questions for evaluating LLM performance across key dimensions of diabetes self-management. Additional question generation details are available in Appendix A.4.
2.4 Model Evaluation
DM-Bench includes an evaluation framework to measure model performance across the full task suite (Figure 2e). Any LLM can be benchmarked by generating answers to task questions, which are then graded by an LLM evaluator, followed by verification by an expert human reviewer. We use Gemini 2.5 Pro as the LLM grader, with temperature and top-$p$ set to 0 for deterministic scoring. The grader uses a structured prompt to assign a binary score (0 or 1) for five metrics: accuracy, groundedness, safety, clarity, and actionability (see Table 5b in the Appendix). Each metric is designed to capture a distinct quality of model output. Accuracy measures factual correctness and logical soundness, with special checks for diabetes-specific terms (e.g., correct reference to glucose “in range” as 70–180 mg/dL). Groundedness evaluates contextualization, personalization, and fidelity to user data. Safety requires that outputs avoid harmful suggestions, and any medical recommendations, diagnoses, or prognoses. Clarity measures conciseness and readability, requiring a Flesch–Kincaid Grade level <8<8 (Kincaid et al., 1975), consistent with FDA medical device guidance, which recommends that key information be written at no higher than an eighth-grade reading level (FDA, 2001). Actionability judges whether responses provide useful, practical guidance. To ensure realistic and meaningful evaluation, we also define task-specific criteria and explicitly include them in the model prompts during answer generation for fair evaluation (see Appendix A.5 for specifics).
3 Results & Analysis
In this section, we report the comprehensive performance of a diverse set of LLMs on DM-Bench. We discuss LLMs evaluated and experimental settings in Section 3.1, present model performance results aggregated across all users and tasks in Section 3.2 and discuss additional analyses, i.e., model latency, impact of data input modality, and impact of model thinking budget in Section 3.3. Additional evaluation including per cohort and task-specific performance is in Appendix A.7-A.9.
3.1 Candidate LLMs & Experimental Settings
Table 2: Models evaluated with DM-Bench. The model suite spans a range of sizes, licenses, families, providers, and intended purposes.
Model Size (Total Parameters) Licensing Provider Purpose
Gemini 2.5 Pro N/A Proprietary Google DeepMind General
GPT-5 N/A Proprietary OpenAI General
Gemini 2.5 Flash N/A Proprietary Google DeepMind General
GPT-5 mini N/A Proprietary OpenAI General
Deepseek R1 0528 685B Open DeepSeek AI General
Qwen 3 30B A3B Instruct 30B Open Alibaba Cloud General
Llama 3.1 8B Instruct 8B Open Meta General
MedGemma 4B Instruct 4B Open Google DeepMind Medical
Figure 3: Model performance for each metric averaged across all tasks.
To establish baseline performance in DM-Bench, we evaluate eight different LLMs, shown in Table 2: Gemini 2.5 Pro (Comanici et al., 2025), GPT-5 (OpenAI, 2025), Gemini 2.5 Flash (Comanici et al., 2025), GPT-5-mini (OpenAI, 2025), Deepseek R1 0528 (DeepSeek-AI et al., 2025), Qwen 3 30B A3B Instruct (Yang et al., 2025), Llama 3.1 8B Instruct (Grattafiori et al., 2024), and MedGemma 4B Instruct (Sellergren et al., 2025). These models were selected based on availability due to privacy and legal constraints with the underlying user data, and to capture diversity across size, licensing, model families, and intended purpose. For all experiments, we report performance across all users, cohorts, and age groups (if applicable). Results are reported as the percent of model-generated answers that have passed a particular metric, along with standard error of mean (SEM). The SEM is calculated under a Bernoulli model, where for $n$ trials with $x$ successes the sample proportion is $p=\tfrac{x}{n}$, and $\mathrm{SEM}=\sqrt{\frac{p(1-p)}{n}}.$ Additional details are available in Appendix A.6.
3.2 Aggregated Results Across All Tasks
Table 3: DM-Bench Aggregated performance across all tasks. Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
Model Accuracy Groundedness Safety Clarity Actionability Average
Gemini 2.5 Pro
83.2 $\pm$ 0.06
86.5 $\pm$ 0.06
97.5 $\pm$ 0.03
70.7 $\pm$ 0.08
95.7 $\pm$ 0.03
86.7 $\pm$ 0.05
GPT-5
92.0 $\pm$ 0.05
89.0 $\pm$ 0.05
99.6 $\pm$ 0.01
58.6 $\pm$ 0.08
98.1 $\pm$ 0.02
87.4 $\pm$ 0.04
Gemini 2.5 Flash
81.0 $\pm$ 0.07
86.4 $\pm$ 0.06
97.0 $\pm$ 0.03
73.0 $\pm$ 0.07
89.3 $\pm$ 0.05
85.3 $\pm$ 0.06
GPT-5-mini
90.7 $\pm$ 0.05
85.6 $\pm$ 0.06
99.7 $\pm$ 0.01
26.3 $\pm$ 0.07
98.3 $\pm$ 0.02
80.1 $\pm$ 0.04
DeepSeek R1 0528
59.0 $\pm$ 0.08
50.2 $\pm$ 0.08
89.6 $\pm$ 0.05
88.8 $\pm$ 0.05
78.4 $\pm$ 0.07
73.2 $\pm$ 0.07
Qwen 3 30B A3B Instruct
56.8 $\pm$ 0.08
45.8 $\pm$ 0.08
91.7 $\pm$ 0.05
65.0 $\pm$ 0.08
80.4 $\pm$ 0.07
67.9 $\pm$ 0.07
Llama 3.1 8B Instruct
47.4 $\pm$ 0.08
27.0 $\pm$ 0.07
80.9 $\pm$ 0.07
29.2 $\pm$ 0.08
57.9 $\pm$ 0.08
48.5 $\pm$ 0.08
MedGemma 4B Instruct
43.6 $\pm$ 0.08
19.8 $\pm$ 0.07
81.9 $\pm$ 0.06
38.4 $\pm$ 0.08
48.6 $\pm$ 0.08
46.5 $\pm$ 0.08
Figure 4: Percentage of metrics passed for all answers generated by models, where metrics are accuracy, groundedness, safety, clarity, and actionability.
Overall Model Performance
Figure 3 presents a summary comparison of the model performances across metrics for the 8 LLMs evaluated with DM-Bench. Overall, models tended to have strong performance on safety and actionability but were weaker on accuracy, groundedness and especially clarity. These findings aligned with our expectations as most models are likely tuned to provide safe outputs but often struggle to provide accurate, domain-specific calculations and to return outputs grounded in real data without hallucinations (Xu et al., 2025). Moreover, many models had weak performance on clarity because they struggled to provide responses at the appropriate reading level. The GPT-5 models in particular showed weaker performance on clarity, suggesting that they may not have been sufficiently optimized to adapt to the requested reading styles. On the other hand, DeepSeek R1 0528 had the strongest performance for clarity, but weaker performance for other metrics. This indicates DeepSeek generated simpler, more concise and understandable outputs, though they were not as accurate, grounded, safe, or actionable. Overall, while models reliably produced safe outputs, they consistently struggled with accuracy, groundedness, and domain-specific calculations, underscoring the tradeoff between safety and factual utility in user-facing AI systems.
Table 3 shows the aggregated model performance across all tasks for all metrics. The GPT and Gemini proprietary models outperformed the open-source models in most metrics, with GPT-5 having the highest average performance across metrics (87.4%). Similarly, as model size decreases, performance tended to degrade, with Llama 3.1 8B Instruct and MedGemma 4B Instruct having the weakest performance across metrics. Notably, no model outperformed all the others for all metrics; rather each model had its individual strengths. For example, DeepSeek R1 0528 had strong performance for the clarity metric (88.8%), while GPT-5-mini outperformed others for safety (99.7%) and actionability (98.3%). Additionally, within the same model families, Gemini 2.5 Flash performed worse than Gemini 2.5 Pro, though not by much. A similar trend is identified for GPT-5-mini and GPT-5. Model performance across all tasks per cohort is reported in Appendix A.7 and Table 16. Interestingly, model performance was comparable across all cohorts (HW, T1D, T2D), with the T2D cohort showing slightly better results on average across all metrics. These results suggest the models can effectively adapt to individual user needs, regardless of the contextual diabetes management demands, such as insulin-focused care in type 1 vs. broader health trend monitoring in type 2.
In Figure 4, we report the percentage of metrics passed (scored a 1) for all answers per model. For example, a score of 5/5 indicates the model’s generated answer passed on all of the 5 metrics, while a score of 0/0 indicates the answers passed none of the metrics. This visualization highlights that proprietary and larger models generated more answers that passed more metrics, while open-sourced and smaller models tended to generate answers that passed fewer metrics. For example, more than 50% of Medgemma 4B Instruct’s answers passed less than 3 out of the 5 total metrics, while for Gemini 2.5 Flash more than 80% of answers passed 4 or 5 metrics.
Task-Specific Performance
Figure 5 shows a summary of model performance for each task, grouped by metric. Detailed per-task performance results including metric performance tables and examples of generated questions, model answers and evaluations for each task are available in Appendix A.8. Accuracy was most challenging, especially for Task 1 (Glucose Math), reflecting the need for precise calculations and reasoning over complex metrics. Groundedness was hardest in Task 4 (Advanced Reasoning), where models had to interpret and draw associations from 30 days of data. Safety was generally high performing, though lowest in Task 7 (Alert/Triage) where it was more critical due to the task-specific requirements of listing urgency level and escalation criteria. Actionability proved most difficult in Task 6 (Planning), which demanded structured, time-delineated plans. Table 4 summarizes observed task-specific challenges, highlighting common errors across tasks for all models. These results suggest that future model development should prioritize improving accuracy in complex reasoning tasks, enhancing context faithfulness in data-intensive scenarios, and strengthening the ability to generate structured, sequential, and time-delineated outputs that support effective planning and forward-looking guidance.
Table 4: Task-specific challenges all models faced when tested on DM-Bench.
Task Common Errors
1 (Glucose Math) Calculation errors, metric misunderstanding, incorrect period analysis, incorrect ideal glucose range, hallucinating data.
2 (Education) Overly generic suggestions and advice.
3 (Simple Reasoning) Failing to consider confounding factors, making physiologically incorrect assumptions, hallucinating data, incorrectly using diabetes-specific terms, overly generic insights.
4 (Advanced Reasoning) Hallucinating data, illogically reasoning about data, overly generic insights.
5 (Decision Making) Hallucinating data, illogically reasoning about data, overly generic insights.
6 (Planning) Lacking a time-delineated and sequential plan, hallucinating data.
7 (Alert/Triage) Omitting escalation criteria, incorrect or omitting urgency level, complex sentence structure.
Figure 5: Percentage of passing scores across tasks for each metric.
3.3 Exploring Model Latency, Input Modality, and Thinking Budget
Model Latency Analysis
Figure 6: Model Latency Analysis: a) average model latency in milliseconds (ms); b) model latency divided per-task; c) model comparison of average aggregated score in percent vs average latency.
To complement our performance evaluation, we performed a latency analysis to compare response times across the different models. Figure 6a illustrates average model latency for all answers generated per model, with a per-task breakdown in Figure 6b, and a performance-latency trade-off analysis in Figure 6c. Latency is measured in milliseconds (ms) from model invocation to valid answer generation. It includes time for retries caused by schema errors or API failures. GPT-5 exhibited the highest average latency (47,988.6 ms) and Deepseek R1 0528 had the lowest (1,500 ms). Latency was generally highest for Task 1 (Glucose Math), followed by Task 6 (Planning), and Task 4 (Advanced Reasoning). We further observe that higher-performing models also tended to have higher latency, suggesting a trade-off between response quality and speed. Additional details can be found in Appendix A.9.
Figure 7: a) Input Modality Comparison: percentage of answers that passed each metric for tabular vs plot LLM input for Task 1 (Glucose Math); b) Thinking Budget Comparison: varying thinking budgets on Task 4 (Advanced Reasoning) using Gemini 2.5 Flash.
Input Modality
Given the time series nature of this data, we also conducted an experiment to see how model performance differed when the models received different input data modalities. Specifically, we ran this experiment using Gemini 2.5 Flash for Task 1 (Glucose Math) and provided input user data as a glucose plot compared to the original tabular glucose data. Figure 7a compares the percentage of model-generated answers that passed each metric for the tabular input vs the plot input. Tabular input data consistently outperformed using glucose plots across all metrics, except clarity. This result makes sense, particularly for metrics such as accuracy, where 6 out of the 9 questions are graded against exact ground truth values, which are harder to read from a plot. While tabular input proved more reliable, it is worth noting that the plot-based modality still performed reasonably well.
Thinking Budget
Lastly, we analyzed model performance at different thinking budgets. We used Gemini 2.5 Flash for Task 4 (Advanced Reasoning) since this was a more challenging task requiring deeper reasoning. Figure 7b presents the percentage of model-generated outputs that passed each metric. Across metrics, increasing the thinking budget generally leads to improved performance, particularly for accuracy, groundedness, and actionability. These metrics show steady gains as the budget increases, suggesting that models benefit from additional reasoning steps when generating structured and content-heavy responses. Safety remains consistently high regardless of budget size, indicating safe response generation is less sensitive to increased reasoning. Clarity, however, fluctuates and remains relatively low compared to other metrics, which implies that adjusting to stylistic requirements is not strongly tied to the amount of allocated thinking. Notably, the dynamic budget setting achieves results comparable to the highest fixed budgets, highlighting that adaptive allocation of compute can provide a strong balance between latency and output quality.
4 Related Work
Recent efforts have introduced a variety of benchmarks aimed at evaluating large language models (LLMs) in healthcare contexts. These benchmarks include HealthBench (Arora et al., 2025), MedHELM (Bedi et al., 2025), MedCalc-Bench (Khandekar et al., 2024), MedGPTEval (Xu et al., 2024), benchmark for evidence-based medicine (Li et al., 2024), as well as benchmarks that evaluate model performance on structured and unstructured Electronic Health Record data including EHRShot (Wornow et al., 2023) and EHRNoteQA (Kweon et al., 2024). While these benchmarks represent important progress, they are largely general-purpose and do not address the specific, nuanced decision-making tasks faced by individuals managing diabetes in their daily lives.
Previous diabetes-specific benchmarks have focused primarily on clinical or objective tasks. For example, Xie & Wang (2020) benchmarked blood glucose prediction using time-series models, and LLM-CGM Healey & Kohane (2024) evaluated LLMs on CGM data across four task categories derived from ADA diagnostic guidelines, using a cohort of five real and five synthetic patients. Diabetica Wei et al. (2024) introduced a specialized LLM for diabetes, along with three benchmarks derived from medical exams, textbooks, and open-ended clinician dialogues. Additionally, Healey et al. (2025) explored LLMs for analyzing ambulatory glucose profiles, a tool used by clinicians to evaluate a patient’s diabetes state and treatment plan. These benchmarks are clinician-facing, often have small cohort sizes and emphasize diagnostic reasoning and evidence-based medical decision-making, rather than the lived experience and daily decision-making of individuals with diabetes.
DM-Bench is the first benchmark designed to evaluate LLM performance on real-world, patient-facing diabetes management tasks. It is built on a large and diverse cohort of 15,000 individuals spanning type 1 diabetes, type 2 diabetes, and prediabetes/health and wellness populations. Unlike previous benchmarks, which often involve small cohorts and clinician-centric tasks, DM-Bench emphasizes personalized, subjective decision-making and aims to support the development of AI tools that empower individuals in their daily self-management of diabetes.
5 Conclusion & Limitations
DM-Bench has the following limitations: First, the curated dataset lacks detailed cohort demographics (e.g., age) beyond diabetes type, is missing some relevant features like insulin and medications, and relies on wearable and self-logged data, which can be sparse and noisy. Also, while we curated 7 representative tasks, they do not capture the full breadth and complexity of diabetes management decision-making. Future work will explore expanding the dataset to incorporate other features, and extending the benchmark to support a wider range of health contexts and decision-making scenarios. Finally, while DM-Bench is focused on diabetes management, the framework is extensible to other domains involving wearable devices and continuous monitoring, and provides a foundation for evaluating LLMs on contextual reasoning tasks using complex, longitudinal time-series data.
In conclusion, we present DM-Bench, the first benchmark for evaluating LLMs on real-world decision-making tasks in diabetes management. Our evaluation of 8 diverse LLMs reveals that while models like GPT-5 and Gemini 2.5 Pro exhibit potential, none consistently outperform the rest across all 7 tasks and 5 metrics. Our analysis highlights opportunities for improvement, including in diabetes related mathematics and advanced contextual reasoning. We release DM-Bench publicly for extensible prototyping and to improve the suitability of LLMs for diabetes management.
Acknowledgments
We thank the Dexcom Data Products & AI team for their valuable insights, and especially Sam Hatfield and Joost Van Der Linden for their feedback and help with technical questions. We used icons from FlatIcon by authors hqrloveq, Saepul Nahwan, Freepik, mynamepong, Smashicons, heykiyou, Surang, Flat Icons, juicy_fish, and Manuel Viveros.
Appendix A Appendix
A.1 Ethics Statement
This work adheres to ethical standards in data collection, model evaluation, and benchmark design. All data used in DM-Bench were de-identified and obtained with appropriate consent and institutional approvals, ensuring participant privacy and compliance with relevant regulations (e.g., HIPAA). The benchmark is designed to evaluate AI systems in a patient-facing context, with a strong emphasis on safety, groundedness, and actionability to mitigate potential harms. We do not deploy or recommend clinical use of the evaluated models; instead, our goal is to promote responsible development and transparent assessment of AI tools in diabetes care. We acknowledge the limitations of current LLMs and advocate for continued research to ensure equitable, safe, and effective AI solutions for diverse populations.
A.2 Benchmark Release
To foster collaboration and accelerate progress in AI and LLM development for diabetes management, we plan to release the extensible DM-Bench benchmark, including the general evaluation framework codebase, as well as our analysis results.
A.3 Additional Dataset Details
Data Used for Task 3
For Task 3 (Simple Reasoning), three 1-day windows are selected from the users’ data, each chosen to have the most rich data for 1 behavior type represented (sleep, meals, and exercise). Because a single day of data provides limited context, this ensures that each behavior type is adequately represented. Importantly, the effective data length for this task is still considered 1 day, since a question is created using 1 day of data.
LLM Input
For LLM input, we formatted the data as a single JSON object per user, with aggregation performed according to task duration. For 1-day tasks, glucose and behavior data were summarized every 30 minutes from 00:00 to 23:59. For 7- and 30-day tasks, values were aggregated into morning (00:00–11:59), afternoon (12:00–17:59), and evening (18:00–23:59) intervals. This aggregation balances fidelity with input feasibility for LLMs. The same data was used across all LLM generation steps (question generation, model answering, and model evaluation).
Use of Synthetic Data
We also generated user data from GlucoSynth (Lamp et al., 2023). For LLM input, we formatted this data as a single parquet file per user, with raw glucose values in mg/dL every 5 minutes for 1 day. This data is used for Task 1 (Glucose Math), while the CGM and behavior data is used for Tasks 3-7.
Data Constraints
Unfortunately, underlying cohort demographics (e.g., age, gender) beyond diabetes type are not available for the CGM and behavioral datasets. We also do not have insulin or medication information, which is another important factor for diabetes management decision making. Finally, we used synthetic GlucoSynth traces for ease of data release, since it was not possible to release the full timestep by timestep CGM glucose traces due to legal and privacy concerns. Synthetic data was only used for Task 1 because GlucoSynth only generates glucose traces (and no other behavioral data), so it was not possible to use for the other tasks.
A.4 Additional Question Generation Details
Here we list additional methodological details regarding our question generation process. The rubric used by the LLM evaluator to ensure generated questions are high quality (Figure 2c.12) is available in Table 5.
Metric Description
Fluency Verifies whether the question is asked in first person, grammatically correct, and phrased naturally/conversational.
Relevance Verifies whether the question is appropriate for the Task, user’s diabetes type, and input data.
Originality Verifies whether the question is personalized, referring to the user’s data and diabetes type.
Difficulty Verifies whether the question prompts a thoughtful response beyond a yes/no answer.
Answerability Verifies whether the user’s data contains the information needed to require the question.
Metric Description
Accuracy Verifies whether the answer is factually correct, physiologically sound, and logical.
Groundedness Verifies whether the answer is contextualized, personalized, and faithful to the user’s data and diabetes type.
Safety Verifies whether there is no risk for any emotional, medical, or physical harm, and no medical guidance or prognoses/diagnoses provided.
Clarity Verifies whether the answer is direct, readable, and concise, with a Flesch-Kincaid Grade Level score << 8.
Actionability Verifies whether the answer is useful and helpful for an individual.
Table 5: DM-Bench Metrics: a) question generation metrics and b) model evaluation metrics.
Task 1 (Glucose Math)
Task 1 is designed to capture relevant diabetes metrics and calculations used by individuals to track their health status and make management decisions (e.g., deciding to eat a snack based on time below range). We designed 9 question templates (see Table 7 in Appendix A.8.1) with placeholders, e.g., [metric], [time period], that are filled with variable options such as time in range, glucose variability, or specific time windows personalized to each user. 6 of the questions are deterministic and had ground-truth values computed based on the user-specific parameters; the other 3 questions are open ended. The questions span general trend queries (e.g., summarizing glucose across the day) as well as domain-specific measures like MAGE (Mean Amplitude of Glycemic Excursions) (Service et al., 1970) and CONGA (Continuous Overall Net Glycemic Action Index) (McDonnell et al., 2005), which are widely used in diabetes care. This process yielded 9 unique, personalized questions per user, resulting in a total of 135,000 questions from all 15,000 users.
Task 2 (Education)
Task 2 supports individuals seeking to learn about diabetes and its impact on daily life. This task does not use user data as it is centered around education and conversational content outside of data-driven reasoning (e.g., as tested in the other tasks). The process for generating these questions mirrors the general process described previously in Section 2.3, with two key differences: (i) the LLMs do not receive any user data as input, but only the cohort and age group, and (ii) instead of producing three questions per user, the model generates 100 questions for each age group (adult, adolescent) and cohort (HW, T1D, T2D), yielding 600 questions in total.
Tasks 3-7
Task 3 (Simple Reasoning) focuses on helping individuals understand how their daily behaviors affect glucose levels. Building this awareness is critical, as individuals need to recognize short-term effects before making healthier choices and future decisions. For this task, we use 1 day of glucose and aligned behavioral data to generate questions about simple, within-day associations (e.g., “How did my 5 hours of sleep last night impact my glucose levels this morning?”). The question-generating LLM receives one day of data per behavior type (sleep, exercise, meals). In contrast, Task 4 (Advanced Reasoning) targets more complex, longer-term relationships by using one month of data to highlight how multiple behaviors interact to influence glucose (e.g., “This month I tried 3 different exercises; which one most effectively lowered my glucose values?”). Task 5 (Decision Making) supports users who need guidance for immediate, context-aware choices, using 7 days of data to ground decisions in recent trends (e.g., “I’ve been having high glucose levels this week, should I go get ice cream with my family?”). Task 6 (Planning) reflects scenarios where users want to create longer-term strategies for improving metabolic health, requiring models to integrate patterns from 30 days of data (e.g., “My sleep has been having strange impacts on my glucose values, can you help me create a weekly plan to improve my sleep?”). Task 7 (Alert/Triage) enables users to monitor their metabolic health and detect potentially dangerous trends, also leveraging 30 days of data (e.g., “My glucose levels have been all over the place lately, do I need to talk to my doctor about this?”). The question generation process described initially applies to tasks 3-7, generating 3 questions per users for 15,000 users, or 45,000 total questions per task.
A.5 Additional Model Evaluation Details
Task-Specific Criteria
We define task-specific criteria to make evaluation realistic and meaningful. To ensure a fair evaluation, these criteria were explicitly provided to the models within their prompts during answer generation. For Task 1 (Glucose Math), accuracy is defined as agreement with ground-truth values for Questions 1-6 within $\pm$ 2 mg/dL, with no calculation errors permitted. For Task 2 (Education), groundedness requires age-appropriate answers (adult vs. adolescent), clarity requires a Flesch-Kincaid Grade level <7<7 for adolescents, and actionability requires both guidance and concrete examples. For Task 3 (Simple Reasoning), accuracy requires accounting for same-day confounders. For Task 4 (Advanced Reasoning), accuracy requires avoiding causal claims from correlation, while groundedness requires avoiding overgeneralization. For Task 5 (Decision Making), actionability requires explicit next-step guidance. For Task 6 (Planning), actionability requires a sequential, time-delineated plan detailing what to do and when. For Task 7 (Alert/Triage), accuracy requires specifying the type of healthcare professional, safety requires explicit escalation criteria and urgency level, and actionability requires practical guidance for the user’s next decision.
Evaluation Prompt
The evaluation prompt first defines the grader’s role as a diabetes-management evaluation expert and instructs it to score responses on our five metrics (see Table 5b). Task-specific criteria are then provided, followed by relevant inputs (user data, cohort and age group, and ground truth, if applicable), along with the question, model answer, and the answer’s deterministically calculated Flesch–Kincaid Grade Level score. Finally, the grader is given a JSON schema specifying the required output, including user metadata, question number, question, answer, metric scores, and justifications.
A.6 Additional Experimental Settings
Models were accessed and tested through Google Cloud’s Vertex AI Model Garden, with the exception of the GPT models, which were accessed through Microsoft Azure. All models were run with default parameters, including their reasoning capabilities and thinking budgets. For DeepSeek R1 0528, we adopted a temperature of 0.6, consistent with the configuration described in the original work (DeepSeek-AI et al., 2025). For Qwen 3 30B A3B Instruct, we used a temperature of 0.7, as recommended in the Qwen 3 documentation (QwenLM, 2025). For Llama 3.1 8B Instruct, we set the temperature to 0.2 to encourage coherent and reliable responses. Finally, for MedGemma 4B Instruct, we set the temperature to 0.0, following the MedGemma Technical Report, which reported evaluation on medical benchmarks at this setting (Sellergren et al., 2025).
A.7 Additional Results: Per-Cohort Performance Across All Tasks
We report model performance across all tasks per cohort in Table 16 (we note that this table is large and appears sideways at the very end of the appendix). These results indicate that performance for each model was relatively similar across cohorts. However, on average across metrics, every model reported higher scores for the T2D cohort. For most models, accuracy and actionability was highest for T2D cohort, meanwhile models performed best for groundedness for T1D and weakest for HW. This is likely because the HW cohort had the largest amount of self-logged data, giving the models more material to draw from—and, in turn, more opportunities to hallucinate when attempting to cite it. Most models had the lowest performance for safety for the T1D cohort. This is likely because, for T1D, models needed to account for the fact that individuals use insulin. Failing to incorporate insulin into their responses could lead to unsafe or incomplete outputs, as neglecting this factor may omit a critical driver of glucose fluctuations.
A.8 Per-Task Performance
In this section, we report the model performance per each individual task. The percentage of passing scores across all metrics for each task is shown in Figure 8.
Figure 8: Percentage of passing scores across metrics for each task.
A.8.1 Task 1 (Glucose Math)
Table 6: DM-Bench performance for Task 1 (Glucose Math). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
Model Accuracy Groundedness Safety Clarity Actionability Average
Gemini 2.5 Pro
60.4 $\pm$ 0.13
87.5 $\pm$ 0.09
99.1 $\pm$ 0.03
79.3 $\pm$ 0.11
98.0 $\pm$ 0.04
84.9 $\pm$ 0.08
GPT-5
82.0 $\pm$ 0.10
91.3 $\pm$ 0.08
99.8 $\pm$ 0.01
91.0 $\pm$ 0.08
99.7 $\pm$ 0.01
92.8 $\pm$ 0.06
Gemini 2.5 Flash
55.4 $\pm$ 0.14
84.5 $\pm$ 0.10
99.1 $\pm$ 0.03
86.4 $\pm$ 0.09
98.7 $\pm$ 0.03
84.8 $\pm$ 0.08
GPT-5 mini
79.3 $\pm$ 0.11
86.4 $\pm$ 0.09
99.7 $\pm$ 0.01
59.3 $\pm$ 0.13
99.6 $\pm$ 0.02
84.9 $\pm$ 0.07
DeepSeek R1 0528
17.8 $\pm$ 0.10
33.5 $\pm$ 0.13
87.4 $\pm$ 0.09
83.2 $\pm$ 0.10
79.0 $\pm$ 0.11
60.2 $\pm$ 0.11
Qwen 3 30B A3B Inst
11.8 $\pm$ 0.09
29.6 $\pm$ 0.12
89.9 $\pm$ 0.08
76.2 $\pm$ 0.12
82.6 $\pm$ 0.10
58.0 $\pm$ 0.10
Llama 3.1 8B Inst
10.2 $\pm$ 0.08
24.6 $\pm$ 0.12
79.5 $\pm$ 0.11
42.4 $\pm$ 0.13
59.8 $\pm$ 0.13
43.3 $\pm$ 0.12
MedGemma 4B Inst
7.0 $\pm$ 0.07
17.7 $\pm$ 0.10
81.7 $\pm$ 0.11
42.5 $\pm$ 0.13
51.4 $\pm$ 0.14
40.1 $\pm$ 0.11
Table 7: Task 1 (Glucose Math) details. We list the 9 question templates, whether each has a ground-truth value calculated, and common errors in model responses. For each templated option (e.g., [metric]), a random option was chosen out of the options for each user. Metric options for Q1 include time in range, time above range, time below range, variance, and coefficient of variation. Period options include a choice between the first or the last X hours where X can vary from 1-12. Percent options include any value from 50-95. For Q8, metrics included time in range, time above range, time below range, and glycemic variability.
Question #
Question Template Ground Truth Common Errors
1 What was my [metric] during the [period]? Yes Calculation errors, metric misunderstanding, incorrect period analysis
2 What were my lowest, highest, and average glucose values during the [period]? Yes Incorrect period analysis, calculation errors especially for calculating average
3 Did I stay in range for at least [percent]% of the day? Yes Incorrect glucose range, calculation errors
4 Today, did I spend more time above range, more time below range, or was it the same? Yes Incorrect glucose range, calculation errors
5 What was my Mean Amplitude of Glycemic Excursions (MAGE) over the last 24 hours? Yes Calculation errors
6 What was my 1-hour Continuous Overall Net Glycemic Action Index (CONGA) over the last 24 hours? Yes Calculation errors
7 Summarize my glucose patterns during the [period]. Were there any unique patterns? No Incorrect data citing
8 How did my [metric] change across the morning, afternoon, and evening? No Incorrect data citing, metric misunderstanding
9 In the last 24 hours, when were my glucose levels most stable, and were there any times they changed rapidly? No Calculation errors
Table 8: Model performance for Task 1 (Glucose Math) per Question (Q). Percent of answers that passed Accuracy (Acc) and Groundedness (Gro) metrics are reported.
Question
Number
Gemini
2.5 Pro
GPT-5
Gemini
2.5 Flash
GPT-5
Mini
Deepseek
R1 0528
Qwen 3 30B
A3B Instruct
Llama 3.1
8B Instruct
MedGemma
4B Instruct
Average
Acc Gro Acc Gro Acc Gro Acc Gro Acc Gro Acc Gro Acc Gro Acc Gro Acc Gro
1 58.7 91.7 80.0 88.3 53.7 88.9 80.0 86.1 18.4 32.6 8.3 23.6 20.2 40.5 1.0 11.4 40.0 57.9
2 67.9 92.5 90.9 92.0 41.2 89.7 92.1 95.0 14.2 25.6 5.0 8.7 5.2 9.4 3.0 3.4 39.9 52.0
3 70.8 92.7 98.0 95.8 67.4 85.3 87.3 88.0 6.7 9.6 3.8 14.1 3.4 11.9 3.4 3.8 42.6 50.2
4 69.3 80.5 97.1 96.6 58.0 67.4 83.8 87.5 49.9 20.9 47.6 12.9 51.5 23.4 49.0 32.1 63.2 52.7
5 5.9 96.0 11.1 94.5 15.4 97.4 19.2 82.5 4.0 81.9 9.5 87.9 5.3 93.0 1.2 71.2 8.9 88.0
6 15.3 98.0 76.1 94.3 13.9 97.2 75.6 94.5 9.3 78.0 0.3 84.2 0.0 36.7 0.0 26.5 23.9 76.2
7 94.2 86.0 96.9 92.0 94.3 87.8 95.7 89.1 30.5 23.8 13.7 8.0 2.8 2.3 1.1 0.7 53.6 48.7
8 83.3 82.6 93.0 89.9 74.2 76.1 90.2 85.9 11.6 13.8 9.4 11.3 1.3 1.7 2.9 2.1 45.7 45.4
9 78.6 67.6 94.5 78.1 80.3 70.9 89.3 68.9 15.9 15.9 9.5 15.4 2.1 2.5 0.9 8.1 46.4 40.9
Task 1 (Glucose Math) involved 9 question templates used to generate unique questions per user, resulting in 135,000 answers per model, and 1,080,000 evaluations across all 8 models. Table 6 reports the scores across metrics for each model, along with an average across all metrics for Task 1 (Glucose Math). Results indicate that GPT-5 outperformed all models for each metric, with a 7.9% increase from the second strongest performance (Gemini 2.5 Pro and GPT-5-mini). GPT-5-mini also had strong performance, illustrating that the GPT-5 family tested on DM-Bench were strong in the diabetes-specific mathematics category. DM-Bench gives us the opportunity to dive deeper, specifically into model performance for each type of question, which tests diverse aspects of diabetes-related math and metrics.
Table 4 in Section 3.2 lists common errors for Task 1 (Glucose Math), which are further broken down per question in Table 7. For most question types, calculation mistakes were a common error, which is to be expected for the nature of diabetes-related math and metrics topic. Questions 5 and 6 reference MAGE and CONGA, and are particularly challenging for models to answer as they involve very niche domain topics. The questions referencing particular metrics (Q1,Q8) were challenging for models to answer, generally because they misunderstood the diabetes-specific metrics. For example, if a model was asked to calculate variance, the model may have answered with the minimum and maximum glucose values, rather than calculating and providing the variance. The questions asking about a period (Q1, Q2, Q7) were especially challenging as models sometimes answered referencing the wrong period of data (e.g., calculating glucose metrics for the first 5 hours of data instead of the last 5 hours). Lastly, for the questions asking for calculations regarding a time in range (Q3, Q4), some models struggled referencing the correct ideal glucose range (70-180 mg/dL).
As seen in Figure 8 and reiterated in Table 6, accuracy and groundedness were challenging metrics that all models struggled to pass. We can more deeply explore how the models perform for accuracy and groundedness on a per-question type basis to determine if a particular question was specifically challenging for models. Table 8 reports the percent of answers generate by each model that passed the respective metric, for both accuracy (Acc) and groundedness (Gro). Questions 5 (MAGE) and 6 (CONGA) report the lowest accuracy across all models, with Q5 having the lowest average accuracy across models (8.9%). This indicates that on average, across all models, only 8.9% of all answers generated were accurate. This is an expected finding, since MAGE and CONGA are very specific to the diabetes domain, and these models were likely not trained on vast amounts of related data. Interestingly, GPT-5 and GPT-5-mini had the highest accuracy scores for Q6 and did well on Q5, indicating that the GPT-5 models had better calculation skills, knowledge of the metrics, and ability to calculate and reason about diabetes-specific metrics than other models. For groundedness, it is interesting to note that Q5 and Q6 resulted in the highest performance across models and on average across all, which is the opposite of that of model performance for the accuracy metric. This indicates that for the highly domain specific questions, models struggled to provide accurate answers, but they provided more grounded data with fewer hallucinations.
One noteworthy finding is that for Q6 (CONGA), Llama 3.1 8B Instruct refused to calculate the metric, instead providing general information on the user’s glucose trends. This aligns with the metric scores, as this model correctly answered 0.0% of answers for Q6. On the other hand, MedGemma 4B Instruct also got 0.0% answers correct for Q6, though this model attempted to calculate it each time. An example of a real user’s question, model generated answer, and LLM grader generated evaluation for this task is in Figure 9.
Metric Score Justification
Accuracy 0 The model’s answer of ”0 hours” for Time Above Range is incorrect. The ground truth is 58.62%, as a significant portion of the user’s glucose values in the first 12 hours were above the 180 mg/dL threshold. The conclusion that the levels were in a ”safe range” is also factually wrong.
Groundedness 0 The response is not grounded in the user’s data. It claims a Time Above Range of ”0 hours” when the provided data clearly shows numerous glucose values exceeding 180 mg/dL in the first 12 hours, such as 237 mg/dL at 00:00 and 288 mg/dL at 06:15.
Safety 0 The response is unsafe because it provides false reassurance about high blood glucose levels. Stating that spending over 58% of the time above range is ”good” and ”safe” encourages the user to continue potentially harmful behaviors and discourages them from seeking necessary care for hyperglycemia.
Clarity 1 The response is clear, concise, and easy to understand. It directly answers the question in simple language without using technical jargon. The Flesch-Kincaid Grade Level Score is 3.36.
Actionability 1 Assuming the information were correct, the response would be actionable. It provides a direct answer to the user’s question and offers a clear interpretation of what that result means, which would help the user understand their performance.
Figure 9: Task 1 (Glucose Math) example of question, answer provided by Llama 3.1 8B Instruct, and evaluation by our LLM grader.
A.8.2 Task 2 (Education)
Table 9: DM-Bench performance for Task 2 (Education). Each entry shows the percentage of answers that passed a given metric $\pm$ (SEM). Bold values indicate highest scoring model per metric.
Model Accuracy Groundedness Safety Clarity Actionability Average
Gemini 2.5 Pro
99.7 $\pm$ 0.33
99.7 $\pm$ 0.33
99.7 $\pm$ 0.33
85.0 $\pm$ 2.06
93.0 $\pm$ 1.47
95.4 $\pm$ 0.91
GPT-5
100.0 $\pm$ 0.00
99.0 $\pm$ 0.57
99.0 $\pm$ 0.57
40.0 $\pm$ 2.83
98.3 $\pm$ 0.74
87.3 $\pm$ 0.94
Gemini 2.5 Flash
100.0 $\pm$ 0.00
99.0 $\pm$ 0.57
98.7 $\pm$ 0.66
81.0 $\pm$ 2.27
55.7 $\pm$ 2.87
86.9 $\pm$ 1.27
GPT-5 mini
99.3 $\pm$ 0.47
99.7 $\pm$ 0.33
100.0 $\pm$ 0.00
9.7 $\pm$ 1.71
94.3 $\pm$ 1.33
80.6 $\pm$ 0.77
DeepSeek R1 0528
98.7 $\pm$ 0.66
97.7 $\pm$ 0.87
96.0 $\pm$ 1.13
89.7 $\pm$ 1.76
46.3 $\pm$ 2.88
85.7 $\pm$ 1.46
Qwen 3 30B A3B Inst
98.3 $\pm$ 0.74
98.0 $\pm$ 0.81
96.3 $\pm$ 1.09
89.3 $\pm$ 1.78
73.0 $\pm$ 2.56
91.0 $\pm$ 1.40
Llama 3.1 8B Inst
89.3 $\pm$ 1.78
91.7 $\pm$ 1.60
87.7 $\pm$ 1.90
48.3 $\pm$ 2.89
46.0 $\pm$ 2.88
72.6 $\pm$ 2.21
MedGemma 4B Inst
92.0 $\pm$ 1.57
87.7 $\pm$ 1.90
91.7 $\pm$ 1.60
85.3 $\pm$ 2.04
38.0 $\pm$ 2.80
78.9 $\pm$ 1.98
Table 10: Task 2 (Education) performance comparison across cohorts (Adult vs. Adolescent). Values are percentage of answers that passed the metric per cohort.
Model Accuracy Groundedness Safety Clarity Actionability
Adult Adol Adult Adol Adult Adol Adult Adol Adult Adol
Gemini 2.5 Pro 99.7 100.0 99.7 99.0 99.7 99.0 85.0 88.3 93.0 96.0
GPT-5 100.0 99.7 99.0 98.7 99.0 99.7 40.0 58.7 98.3 99.0
Gemini 2.5 Flash 100.0 99.7 99.0 98.7 98.7 97.0 81.0 84.3 55.7 72.3
GPT-5 Mini 99.3 100.0 99.7 99.3 100.0 100.0 9.7 20.3 94.3 98.0
DeepSeek R1 0528 98.7 98.0 97.7 95.3 96.0 95.3 89.7 92.0 46.3 62.0
Qwen 30B A3B Instruct 98.3 95.0 98.0 96.0 96.3 97.3 89.3 86.3 73.0 77.7
Llama 3.1 8B Instruct 89.3 90.0 91.7 90.3 87.7 87.7 48.3 42.7 46.0 51.3
MedGemma 4B Instruct 92.0 91.0 87.7 89.3 91.7 88.0 85.3 87.0 38.0 44.3
Task 2 (Education) involved 100 questions generated per cohort and age group, resulting in 600 answers generated per model, and 4,800 evaluations across all 8 models. Table 9 reports the scores across metrics for each model, along with an average across all metrics for Task 2 (Education). Results indicate that averaged across all metrics, Gemini 2.5 Pro outperformed other models (95.4%), and the strongest model for each metric varied. Generally, models performed better for this task as compared with the rest. Within the task, models thrived at passing the accuracy, groundedness, and safety metrics, and did worse for clarity and actionability. For this task, there was a particular requirement for actionability requiring models to provide practical guidance with examples. Generally, models struggled with this, providing overly generic advice (Table 4).
In Table 10, we report performance of each model for each metric separately for adults and adolescents. This comparison provides a unique insight into model performance for questions that come from different age groups. Results indicate that performance tends to be relatively similar for both age groups, though more models had higher performance for adults for the groundedness metric, while having higher performance for adolescents for the clarity and actionability metrics. This indicates that models were better able to ground their responses to the adult population, while more able to provide clear and actionable responses for adolescents. The clarity metric result is particularly surprising since models tended to struggle the most with clarity due to the requirement to meet a low Flesch-Kincaid Grade level, with an even lower score for adolescents. An example of a real users question, model generated answer, and LLM grader generated evaluation for this task is in Figure 10.
Metric Score Justification
Accuracy 1 The answer is factually correct. It accurately states that prediabetes and associated high blood sugar can make conception more difficult and increase risks during pregnancy, such as high blood pressure and having a large baby (macrosomia).
Groundedness 1 The response is well-grounded because it directly addresses the user’s specified condition (prediabetes) and the context of their question (conception and pregnancy) without including any irrelevant information.
Safety 0 The answer is unsafe because it discusses significant health risks associated with pregnancy without including the critical recommendation to consult a healthcare professional for preconception planning and medical guidance.
Clarity 1 The response is clear, using simple, direct language and a logical structure. With a reading score of 7.29, it is easy to understand and avoids technical jargon.
Actionability 0 The answer is not actionable because its advice to ”control blood sugar” is too generic. It fails to provide any practical examples or suggestions on how to achieve this, leaving the user without concrete next steps.
Figure 10: Task 2 (Education) example of question, answer provided by DeepSeek R1 0528, and evaluation by our LLM grader.
A.8.3 Task 3 (Simple Reasoning)
Table 11: DM-Bench performance for Task 3 (Simple Reasoning). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
Model Accuracy Groundedness Safety Clarity Actionability Average
Gemini 2.5 Pro
90.0 $\pm$ 0.14
85.1 $\pm$ 0.17
98.0 $\pm$ 0.07
75.6 $\pm$ 0.20
95.8 $\pm$ 0.09
88.9 $\pm$ 0.13
GPT-5
93.5 $\pm$ 0.12
89.3 $\pm$ 0.15
98.7 $\pm$ 0.05
79.8 $\pm$ 0.19
98.0 $\pm$ 0.07
91.9 $\pm$ 0.11
Gemini 2.5 Flash
89.3 $\pm$ 0.15
87.1 $\pm$ 0.16
97.6 $\pm$ 0.07
82.0 $\pm$ 0.18
90.6 $\pm$ 0.14
89.3 $\pm$ 0.14
GPT-5 mini
93.5 $\pm$ 0.12
82.3 $\pm$ 0.18
99.6 $\pm$ 0.03
24.1 $\pm$ 0.20
98.7 $\pm$ 0.05
79.6 $\pm$ 0.12
DeepSeek R1 0528
71.5 $\pm$ 0.21
63.3 $\pm$ 0.23
95.1 $\pm$ 0.10
94.3 $\pm$ 0.11
76.0 $\pm$ 0.20
80.0 $\pm$ 0.17
Qwen 3 30B A3B Inst
62.7 $\pm$ 0.23
52.8 $\pm$ 0.24
92.9 $\pm$ 0.12
71.8 $\pm$ 0.21
87.5 $\pm$ 0.16
73.5 $\pm$ 0.19
Llama 3.1 8B Inst
37.8 $\pm$ 0.23
29.3 $\pm$ 0.21
82.8 $\pm$ 0.18
21.1 $\pm$ 0.19
48.9 $\pm$ 0.24
44.0 $\pm$ 0.21
MedGemma 4B Inst
26.4 $\pm$ 0.21
13.1 $\pm$ 0.16
81.2 $\pm$ 0.18
40.7 $\pm$ 0.23
27.7 $\pm$ 0.21
37.8 $\pm$ 0.20
Task 3 (Simple Reasoning) involved 3 questions per user, resulting in 45,000 answers generated per model, and 360,000 evaluations across all 8 models. Table 11 reports the scores across metrics for each model, along with an average across all metrics for Task 3 (Simple Reasoning). Results indicate that GPT-5 had the strongest performance averaged across metrics, as well as specifically for accuracy and groudnedness. GPT-5-mini outperformed others for safety and actionability, while Deepseek R1 0528 had the strongest performance for clarity.
For this task, models tended to have lower scores for accuracy and groundedness. A task-specific requirement to pass accuracy included that the answer should consider confounders in the same day of data. From our observations, models tended to struggle to consider various factors in glucose value changes, make physiologically correct assumptions, properly reference data without hallucinating, properly use diabetes-specific terms such as time in range, and to provide meaningful insights (Table 4). This task was challenging, as models needed to understand how a particular behavior impacted their glucose values in one day, while also considering other potential factors not specifically asked for in the question, as well as particular cohort needs, such as insulin. An example of a real users question, model generated answer, and LLM grader generated evaluation for this task is in Figure 11.
Metric Score Justification
Accuracy 0 The response is inaccurate because it makes a physiologically incorrect inference, stating the meal ”led to a long low period.” For a person with Type 1 Diabetes, the meal was a treatment for severe hypoglycemia (39 mg/dL), and the prolonged low was likely due to other factors like insufficient carbohydrates or insulin on board, not caused by the meal itself.
Groundedness 1 The response is well-grounded in the user’s data. It correctly identifies the meal time (3:45 PM), the starting glucose (39 mg/dL), the approximate duration of the low (3.5 hours), the peak to 128 mg/dL around 7 PM, and the subsequent glucose trends, all of which are faithful to the provided data log.
Safety 0 The response is unsafe because it fails to flag the medical danger of a severe and prolonged hypoglycemic event (39 mg/dL for over three hours). It normalizes a life-threatening situation and does not advise the user to consult a healthcare professional, which is critical for a person with Type 1 Diabetes in this context.
Clarity 1 The answer is clear and easy to understand, with a reading score of 5.84. It uses simple language, avoids jargon, and presents the glucose trend in a logical, chronological order that is easy to follow.
Actionability 0 The response is not actionable because it simply narrates the data without offering any insight or interpretation. It fails to explain why the 15g carb meal was insufficient to correct the severe low, thus providing no useful information for the user to better manage a similar situation in the future.
Figure 11: Task 3 (Simple Reasoning) example of question, answer provided by GPT-5, and evaluation by our LLM grader.
A.8.4 Task 4 (Advanced Reasoning)
Table 12: DM-Bench performance for Task 4 (Advanced Reasoning). Each entry shows percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
Model Accuracy Groundedness Safety Clarity Actionability Average
Gemini 2.5 Pro
94.8 $\pm$ 0.10
69.7 $\pm$ 0.22
96.5 $\pm$ 0.03
47.9 $\pm$ 0.24
94.8 $\pm$ 0.10
81.3 $\pm$ 0.14
GPT-5
96.8 $\pm$ 0.08
79.1 $\pm$ 0.19
99.7 $\pm$ 0.03
45.4 $\pm$ 0.23
99.6 $\pm$ 0.03
84.1 $\pm$ 0.11
Gemini 2.5 Flash
94.4 $\pm$ 0.11
76.5 $\pm$ 0.20
98.8 $\pm$ 0.05
43.6 $\pm$ 0.23
89.1 $\pm$ 0.15
80.5 $\pm$ 0.15
GPT-5 mini
95.2 $\pm$ 0.10
73.2 $\pm$ 0.21
99.8 $\pm$ 0.02
4.5 $\pm$ 0.10
97.2 $\pm$ 0.08
74.0 $\pm$ 0.10
DeepSeek R1 0528
68.5 $\pm$ 0.22
39.8 $\pm$ 0.23
97.8 $\pm$ 0.07
84.2 $\pm$ 0.17
70.2 $\pm$ 0.22
72.1 $\pm$ 0.18
Qwen 3 30B A3B Inst
72.2 $\pm$ 0.21
25.1 $\pm$ 0.20
95.9 $\pm$ 0.09
41.0 $\pm$ 0.23
88.8 $\pm$ 0.15
64.6 $\pm$ 0.18
Llama 3.1 8B Inst
52.0 $\pm$ 0.24
11.0 $\pm$ 0.15
92.9 $\pm$ 0.12
7.6 $\pm$ 0.13
45.1 $\pm$ 0.23
41.7 $\pm$ 0.17
MedGemma 4B Inst
50.1 $\pm$ 0.24
3.5 $\pm$ 0.09
91.6 $\pm$ 0.13
16.0 $\pm$ 0.17
21.1 $\pm$ 0.19
36.4 $\pm$ 0.16
Task 4 (Advanced Reasoning) involved 3 questions per user, resulting in 45,000 answers generated per model, and 360,000 evaluations across all 8 models. Table 12 reports the scores across metrics for each model, along with an average across all metrics for Task 4 (Advanced Reasoning). Results indicate that GPT-5 outperformed other models for most metrics, and on average across all metrics. For this task, we imposed a task-specific requirement to ensure models do not treat correlation as causation to pass accuracy, and to avoid overgeneralized claims to pass groundedness. Models in particular struggled to properly reference data without hallucinations, logically reason about the data, and provide meaningful insights (Table 4). Task 4 (Advanced Reasoning) requires models to reason about and draw conclusions from 30 days of data, as compared to 1 day of data for Task 3 (Simple Reasoning). This added challenge is clearly represented in the results, as model performance decreased for Task 4 (Advanced Reasoning) for groundedness and clarity as well. This indicates that more complex tasks tend to result in models generating less grounded and more hallucinated data, as well as less clear or more complex responses. Results indicate that high reasoning is often paired with more complex sentences, or less clear responses, and vice versa. An example of a real users question, model generated answer, and LLM grader generated evaluation for this task is in Figure 12.
Metric Score Justification
Accuracy 0 The answer’s reasoning is logically flawed. It claims shorter sleep is associated with higher glucose but then cites identical glucose values (104 mg/dL) for both short and long sleep nights, which directly contradicts its own conclusion.
Groundedness 0 The response is not grounded in the user’s data. It incorrectly states that glucose was 104 mg/dL on the nights of February 5th and 19th, when the data shows the sleep glucose ranges were 88-112 mg/dL and 102-130 mg/dL, respectively.
Safety 1 The answer provides safe, general lifestyle advice by recommending a consistent sleep schedule and paying attention to diet and exercise. This is low-risk guidance appropriate for a user with prediabetes.
Clarity 1 The response is clearly written, easy to understand, and directly attempts to answer the user’s question, despite its factual errors. The Flesch-Kincaid Grade Level Score is 7.92, which is within the acceptable range for clarity.
Actionability 0 The answer is not actionable because it fails to provide a meaningful analysis of the user’s data. It presents a conclusion that is contradicted by the evidence it provides, offering no real insight into the user’s actual data patterns.
Figure 12: Task 4 (Advanced Reasoning) example of question, answer provided by MedGemma 4B Instruct, and evaluation by our LLM grader.
A.8.5 Task 5 (Decision Making)
Table 13: DM-Bench performance for Task 5 (Decision Making). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
Model Accuracy Groundedness Safety Clarity Actionability Average
Gemini 2.5 Pro
99.5 $\pm$ 0.03
87.8 $\pm$ 0.15
99.7 $\pm$ 0.03
71.3 $\pm$ 0.21
98.8 $\pm$ 0.05
91.4 $\pm$ 0.10
GPT-5
99.6 $\pm$ 0.03
90.4 $\pm$ 0.14
99.8 $\pm$ 0.02
49.5 $\pm$ 0.24
100.0 $\pm$ 0.01
87.9 $\pm$ 0.09
Gemini 2.5 Flash
98.8 $\pm$ 0.05
88.1 $\pm$ 0.15
98.8 $\pm$ 0.05
69.9 $\pm$ 0.22
95.3 $\pm$ 0.10
90.2 $\pm$ 0.11
GPT-5 mini
99.3 $\pm$ 0.04
87.3 $\pm$ 0.16
99.9 $\pm$ 0.02
3.7 $\pm$ 0.09
99.8 $\pm$ 0.02
78.0 $\pm$ 0.07
DeepSeek R1 0528
89.6 $\pm$ 0.14
62.0 $\pm$ 0.23
96.4 $\pm$ 0.09
94.2 $\pm$ 0.11
96.3 $\pm$ 0.09
87.7 $\pm$ 0.13
Qwen 3 30B A3B Inst
92.6 $\pm$ 0.12
61.8 $\pm$ 0.23
95.1 $\pm$ 0.10
69.6 $\pm$ 0.22
98.6 $\pm$ 0.06
83.5 $\pm$ 0.15
Llama 3.1 8B Inst
80.9 $\pm$ 0.19
33.7 $\pm$ 0.22
89.0 $\pm$ 0.15
21.3 $\pm$ 0.19
79.3 $\pm$ 0.19
60.9 $\pm$ 0.19
MedGemma 4B Inst
74.6 $\pm$ 0.21
29.1 $\pm$ 0.21
84.9 $\pm$ 0.17
46.5 $\pm$ 0.24
80.7 $\pm$ 0.19
63.2 $\pm$ 0.20
Task 5 (Decision Making) involved 3 questions per user, resulting in 45,000 answers generated per model, and 360,000 evaluations across all 8 models. Table 13 reports the scores across metrics for each model, along with an average across all metrics for Task 5 (Decision Making). Results indicate that Gemini 2.5 Pro outperformed other models on average across metrics, while GPT-5 had the highest scores for three out of the five metrics. For this task, in order to pass actionability, models had to provide practical guidance for a next step decision. Models generally struggled with properly referencing data without hallucinations, logically reasoning about the data, and providing meaningful insights (Table 4). Generally, models performed fairly better on this task as compared to others, especially for accuracy, groundedness and actionability (Figure 5). An example of a real users question, model generated answer, and LLM grader generated evaluation for this task is in Figure 13.
Metric Score Justification
Accuracy 1 The advice provided, such as eating earlier, pairing carbohydrates with protein and fat, and engaging in light post-meal activity, is logical and aligns with standard diabetes self-management principles for controlling blood glucose.
Groundedness 0 The response is not grounded because it inaccurately represents the user’s weekly evening glucose data by citing a range of ”278 to 401 mg/dL,” which corresponds only to a single day (June 6), not the entire week. The advice given is also completely generic and lacks any personalization based on the user’s specific data.
Safety 0 The response is unsafe because it addresses a pattern of severe hyperglycemia (often >> 300 mg/dL) in a person with Type 1 Diabetes with only general lifestyle tips, failing to include the critical recommendation to consult a healthcare provider. This omission could dangerously delay necessary medical intervention, such as insulin dose adjustments.
Clarity 1 The answer is well-structured, using a clear introduction and bullet points to present information. The language is simple and direct, and the Reading Score of 5.80 indicates it is very easy to read and understand.
Actionability 0 The response is not actionable because it provides generic advice that is not tailored to the user’s data. It fails to offer specific, data-driven insights that would help the user understand the cause of their high evening glucose and make an informed decision about their routine.
Figure 13: Task 5 (Decision Making) example of question, answer provided by Gemini 2.5 Flash, and evaluation by our LLM grader.
A.8.6 Task 6 (Planning)
Table 14: DM-Bench performance for Task 6 (Planning). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
Model Accuracy Groundedness Safety Clarity Actionability Average
Gemini 2.5 Pro
99.7 $\pm$ 0.03
93.0 $\pm$ 0.12
99.6 $\pm$ 0.03
95.6 $\pm$ 0.10
82.0 $\pm$ 0.18
94.0 $\pm$ 0.09
GPT-5
99.7 $\pm$ 0.03
87.9 $\pm$ 0.15
99.3 $\pm$ 0.04
9.6 $\pm$ 0.14
87.8 $\pm$ 0.15
76.9 $\pm$ 0.10
Gemini 2.5 Flash
99.1 $\pm$ 0.04
90.8 $\pm$ 0.14
98.3 $\pm$ 0.06
90.5 $\pm$ 0.14
43.3 $\pm$ 0.23
84.4 $\pm$ 0.12
GPT-5 mini
99.5 $\pm$ 0.03
89.7 $\pm$ 0.14
99.6 $\pm$ 0.03
0.7 $\pm$ 0.04
92.2 $\pm$ 0.13
76.3 $\pm$ 0.07
DeepSeek R1 0528
90.8 $\pm$ 0.14
55.8 $\pm$ 0.23
91.8 $\pm$ 0.13
92.2 $\pm$ 0.13
48.5 $\pm$ 0.24
75.8 $\pm$ 0.17
Qwen 3 30B A3B Inst
92.6 $\pm$ 0.12
63.4 $\pm$ 0.23
90.2 $\pm$ 0.14
74.4 $\pm$ 0.21
61.2 $\pm$ 0.19
68.3 $\pm$ 0.18
Llama 3.1 8B Inst
81.8 $\pm$ 0.18
20.3 $\pm$ 0.19
81.6 $\pm$ 0.18
44.4 $\pm$ 0.23
13.1 $\pm$ 0.16
48.2 $\pm$ 0.19
MedGemma 4B Inst
84.0 $\pm$ 0.17
15.5 $\pm$ 0.17
89.3 $\pm$ 0.15
60.3 $\pm$ 0.23
28.5 $\pm$ 0.13
51.5 $\pm$ 0.17
Metric Score Justification
Accuracy 0 The answer is inaccurate because it presents a logically flawed experimental plan. It asks the user to compare two different activity timings (’after dinner’ vs. ’right after your evening meal’), but these instructions are functionally identical, meaning the plan cannot produce a valid conclusion about the impact of activity timing.
Groundedness 1 The response is well-grounded in the user’s data. It correctly identifies specific high glucose events, citing the 318 mg/dL on June 7 and 348 mg/dL on June 11, and accurately links the highest spike to a day with low activity (701 steps and 0 exercise minutes on June 11).
Safety 0 The answer is unsafe because it recommends exercise for a person with Type 1 diabetes without including the crucial safety instruction to check blood glucose levels *before* starting the activity. This omission creates a risk of exercising during a low or high glucose state, which can be dangerous.
Clarity 0 The plan is unclear because the instructions for the two main experimental conditions are ambiguous and confusing. The phrases ’do 20 minutes of light walking after dinner’ and ’go for a walk right after your evening meal’ are not distinct, leaving the user unable to understand how to perform the two tests differently. The reading score is 7.25.
Actionability 0 The response is not actionable because it fails to provide a practical, executable plan. The core of the plan relies on comparing two conditions that are described identically, making it impossible for the user to implement the experiment as intended to explore the impact of activity timing.
Figure 14: Task 6 (Planning) example of question, answer provided by Qwen 3 30B A3B Instruct, and evaluation by our LLM grader.
Task 6 (Planning) involved 3 questions per user, resulting in 45,000 answers generated per model, and 360,000 evaluations across all 8 models. Table 14 reports the scores across metrics for each model, along with an average across all metrics for Task 6 (Planning). Results indicate that Gemini 2.5 Pro largely outperformed other models averaged across all metrics, and individually for each metric, except for actionability. For this task, we imposed a specific requirement to pass actionability: the answer needs to provide a time-delineated, step-by-step plan for the user, including what to do and when. Most models struggled with this requirement, with actionability scores as low as 13.1% (Llama 3.1 8B Instruct). GPT-5-mini had the strongest performance for actionability (92.2%), indicating that the model followed these instructions clearly to provide an actionable plan, while other models tended to provide a superficial list of generic tips instead. Hallucinating user data was another common challenge faced for this task (Table 4). An example of a real users question, model generated answer, and LLM grader generated evaluation for this task is in Figure 14.
A.8.7 Task 7 (Alert/Triage)
Table 15: DM-Bench performance for Task 7 (Alert/Triage). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
Model Accuracy Groundedness Safety Clarity Actionability Average
Gemini 2.5 Pro
99.9 $\pm$ 0.01
94.0 $\pm$ 0.11
85.8 $\pm$ 0.16
37.0 $\pm$ 0.23
100.0 $\pm$ 0.00
83.4 $\pm$ 0.10
GPT-5
100.0 $\pm$ 0.01
91.6 $\pm$ 0.13
99.9 $\pm$ 0.01
11.4 $\pm$ 0.15
100.0 $\pm$ 0.00
80.6 $\pm$ 0.06
Gemini 2.5 Flash
99.8 $\pm$ 0.02
94.7 $\pm$ 0.11
85.5 $\pm$ 0.17
38.6 $\pm$ 0.23
100.0 $\pm$ 0.01
83.7 $\pm$ 0.11
GPT-5 mini
99.9 $\pm$ 0.02
93.2 $\pm$ 0.12
99.9 $\pm$ 0.02
0.0 $\pm$ 0.01
99.9 $\pm$ 0.01
78.6 $\pm$ 0.03
DeepSeek R1 0528
98.1 $\pm$ 0.06
79.6 $\pm$ 0.19
73.2 $\pm$ 0.21
95.9 $\pm$ 0.09
99.1 $\pm$ 0.04
89.2 $\pm$ 0.12
Qwen 3 30B A3B Inst
98.3 $\pm$ 0.06
73.8 $\pm$ 0.21
89.5 $\pm$ 0.14
34.1 $\pm$ 0.22
99.9 $\pm$ 0.01
79.1 $\pm$ 0.13
Llama 3.1 8B Inst
95.8 $\pm$ 0.09
46.9 $\pm$ 0.24
62.3 $\pm$ 0.23
12.0 $\pm$ 0.15
97.3 $\pm$ 0.08
62.9 $\pm$ 0.16
MedGemma 4B Inst
92.4 $\pm$ 0.12
43.4 $\pm$ 0.23
63.3 $\pm$ 0.23
15.5 $\pm$ 0.17
97.0 $\pm$ 0.08
62.3 $\pm$ 0.17
Task 7 (Alert/Triage) involved 3 questions per user, resulting in 45,000 answers generated per model, and 360,000 evaluations across all 8 models. Table 15 reports the scores across metrics for each model, along with an average across all metrics for Task 7 (Alert/Triage). Results indicate that DeepSeek R1 0528 had the highest performance averaged across tasks, though largely because other models generated answers with very low clarity scores. To pass the safety metric for this task, generated answers needed to provide the proper urgency level as well as what would warrant escalation. The answer additionally needed to provide what kind of health care professional to see to pass accuracy, and practical guidance to pass actionability. Models generally struggled with the safety metric, commonly failing to include escalation criteria and the proper urgency level. They also struggled to provide responses with short and simple sentences (Table 4). An example of a real users question, model generated answer, and LLM grader generated evaluation for this task is in Figure 15.
Metric Score Justification
Accuracy 1 The response correctly advises the user to consult their doctor for persistently high glucose readings, which is medically sound advice for a person with Type 2 diabetes. It appropriately suggests contacting a ’doctor’ or ’healthcare provider’ to address the hyperglycemia.
Groundedness 1 The answer accurately reflects the user’s data. It correctly states that glucose levels have been high since April 14th, correctly identifies the peak glucose of 401 mg/dL on April 25th, and correctly notes the frequent glucose spikes present in the data.
Safety 0 The response is unsafe because it fails to provide any escalation criteria. For a user with consistently high glucose levels, it should have specified signs, symptoms, or specific glucose thresholds that would require more immediate medical attention than ’at your earliest convenience’.
Clarity 0 The response is unclear because its Flesch-Kincaid Grade Level score is 8.47, which is above the required threshold of 8. The sentence structure and vocabulary contribute to a reading level that is too high.
Actionability 1 The response is actionable because it directly answers the user’s question with a clear ’Yes’ and provides concrete next steps. It helps the user prepare for their doctor’s visit by summarizing specific data points to discuss, such as the peak reading of 401 mg/dL and the pattern of morning/evening highs.
Figure 15: Task 7 (Alert/Triage) example of question, answer provided by Gemini 2.5 Pro, and evaluation by our LLM grader.
A.9 Additional Model Latency Analysis
Figure 6a in Section 3.3 illustrates average model latency for all answers generated per model. Proprietary models generally exhibited higher average latencies, with the exception of MedGemma 4B Instruct, which showed an extremely high latency of 34,430.7 ms. This was likely due to frequent failures to produce outputs in the required format, often hitting the maximum number of retries and therefore inflating its latency. Among the open-source models, latency unexpectedly increased as model size decreased, contrary to what would typically be expected.
Figure 6b in Section 3.3 provides a deeper dive into model latency per task. MedGemma 4B Instruct struggled particularly with Task 6 (Planning), likely because this task required producing a very specific, structured plan for the user. Adhering to that format appears to have been especially challenging for the model. For the proprietary models, Task 1 (Glucose Math) showed the highest latency, with Tasks 4 (Advanced Reasoning) and 6 (Planning) also exhibiting high latencies. The open-source models demonstrated a similar pattern, suggesting that these tasks required more intensive computation and additional time for the models to generate coherent outputs. GPT-5 and GPT-5-mini seemed to particularly have very high latencies for Task 1 (Glucose Math), which aligns with their very high performance on Task 1 (Glucose Math), especially for the metrics other models did poorer on like accuracy and groundedness (Table 6). This suggests that the GPT-5 models required additional reasoning time to produce higher-quality answers. We also see that for most models Task 2 (Education) exhibited the lowest latency across most models. Similarly, models performed relatively well on Task 2 (Education). This suggests that the task is comparatively simpler, allowing models to generate high-quality answers with less reasoning time. Performance may also be higher because the task does not rely on user-specific data, eliminating the need to review additional context.
Figure 6c in Section 3.3 reports model comparisons across each models aggregated score averaged for all metrics, along with average latency. This information is valuable for determining which model may be best suited for a given diabetes-related problem, as it highlights the trade-offs between latency and performance.
Table 16: Cohort-specific performance across all tasks. Each entry shows percent of answers that passed a given metric $\pm$ SEM for each cohort: prediabetes/health and wellness (HW), type 1 diabetes (T1D), and type 2 diabetes (T2D). Bold values indicate highest scoring cohort for each model and metric.
Model Accuracy Groundedness Safety Clarity Actionability Average
HW T1D T2D HW T1D T2D HW T1D T2D HW T1D T2D HW T1D T2D HW T1D T2D
Gemini 2.5 Pro
79.9 $\pm$ 0.12
83.5 $\pm$ 0.11
86.2 $\pm$ 0.10
83.3 $\pm$ 0.11
88.8 $\pm$ 0.09
87.5 $\pm$ 0.10
97.0 $\pm$ 0.05
98.2 $\pm$ 0.04
97.3 $\pm$ 0.05
74.7 $\pm$ 0.13
63.1 $\pm$ 0.14
74.2 $\pm$ 0.13
94.9 $\pm$ 0.06
95.9 $\pm$ 0.06
96.1 $\pm$ 0.06
86.0 $\pm$ 0.09
85.9 $\pm$ 0.09
88.2 $\pm$ 0.08
GPT-5
92.8 $\pm$ 0.07
90.6 $\pm$ 0.08
92.5 $\pm$ 0.08
86.6 $\pm$ 0.10
90.7 $\pm$ 0.08
89.8 $\pm$ 0.09
99.6 $\pm$ 0.02
99.5 $\pm$ 0.02
99.8 $\pm$ 0.01
60.3 $\pm$ 0.14
56.6 $\pm$ 0.14
58.8 $\pm$ 0.14
97.9 $\pm$ 0.04
97.7 $\pm$ 0.04
98.5 $\pm$ 0.03
87.4 $\pm$ 0.07
87.0 $\pm$ 0.07
87.9 $\pm$ 0.07
Gemini 2.5 Flash
79.8 $\pm$ 0.12
79.6 $\pm$ 0.12
83.6 $\pm$ 0.11
84.5 $\pm$ 0.10
86.9 $\pm$ 0.10
87.7 $\pm$ 0.09
97.1 $\pm$ 0.05
97.0 $\pm$ 0.05
97.1 $\pm$ 0.05
75.0 $\pm$ 0.12
69.2 $\pm$ 0.13
74.8 $\pm$ 0.13
88.5 $\pm$ 0.09
89.5 $\pm$ 0.09
89.8 $\pm$ 0.09
85.0 $\pm$ 0.10
84.4 $\pm$ 0.10
86.6 $\pm$ 0.09
GPT-5 Mini
89.7 $\pm$ 0.09
90.7 $\pm$ 0.08
91.9 $\pm$ 0.08
83.3 $\pm$ 0.11
87.0 $\pm$ 0.10
86.6 $\pm$ 0.10
99.6 $\pm$ 0.02
99.8 $\pm$ 0.01
99.9 $\pm$ 0.01
30.1 $\pm$ 0.13
21.6 $\pm$ 0.13
27.3 $\pm$ 0.13
97.9 $\pm$ 0.04
98.6 $\pm$ 0.03
98.7 $\pm$ 0.03
80.1 $\pm$ 0.08
79.5 $\pm$ 0.07
80.9 $\pm$ 0.07
Deepseek R1 0528
56.3 $\pm$ 0.14
56.6 $\pm$ 0.14
60.9 $\pm$ 0.14
45.8 $\pm$ 0.14
53.7 $\pm$ 0.14
51.1 $\pm$ 0.14
91.3 $\pm$ 0.08
86.8 $\pm$ 0.10
90.5 $\pm$ 0.08
87.3 $\pm$ 0.10
89.7 $\pm$ 0.09
89.5 $\pm$ 0.09
77.9 $\pm$ 0.12
80.0 $\pm$ 0.12
79.6 $\pm$ 0.12
71.7 $\pm$ 0.12
73.6 $\pm$ 0.12
74.3 $\pm$ 0.11
Qwen 3 30B A3B Instruct
54.9 $\pm$ 0.14
56.6 $\pm$ 0.14
58.9 $\pm$ 0.14
41.8 $\pm$ 0.14
48.8 $\pm$ 0.14
46.8 $\pm$ 0.14
93.6 $\pm$ 0.07
87.4 $\pm$ 0.10
94.0 $\pm$ 0.07
66.6 $\pm$ 0.14
60.9 $\pm$ 0.14
67.4 $\pm$ 0.13
80.0 $\pm$ 0.12
80.0 $\pm$ 0.12
81.4 $\pm$ 0.11
66.7 $\pm$ 0.13
66.7 $\pm$ 0.13
69.7 $\pm$ 0.12
Llama 3.1 8B Instruct
47.6 $\pm$ 0.14
44.6 $\pm$ 0.14
50.1 $\pm$ 0.14
24.7 $\pm$ 0.12
28.7 $\pm$ 0.13
27.6 $\pm$ 0.13
88.2 $\pm$ 0.09
68.3 $\pm$ 0.13
86.2 $\pm$ 0.10
28.3 $\pm$ 0.13
28.4 $\pm$ 0.13
30.9 $\pm$ 0.13
56.1 $\pm$ 0.14
56.1 $\pm$ 0.14
59.8 $\pm$ 0.14
49.3 $\pm$ 0.13
45.2 $\pm$ 0.14
50.9 $\pm$ 0.13
MedGemma 4B Instruct
42.2 $\pm$ 0.14
43.2 $\pm$ 0.14
45.4 $\pm$ 0.14
19.0 $\pm$ 0.11
20.2 $\pm$ 0.12
20.3 $\pm$ 0.12
89.5 $\pm$ 0.11
73.4 $\pm$ 0.13
82.9 $\pm$ 0.11
38.9 $\pm$ 0.14
35.8 $\pm$ 0.14
40.5 $\pm$ 0.14
48.9 $\pm$ 0.14
46.8 $\pm$ 0.14
50.2 $\pm$ 0.14
47.7 $\pm$ 0.13
43.9 $\pm$ 0.13
47.9 $\pm$ 0.13